Find the mass of a mineral that has a known density of 2.7 g/cm3and a volume of 14 cm3.
Answer #1

density = 2-79/cm ; volume = 14 cm3 density = masse d = m = m= dxv mass = {2.13/cm) (14 cm) Mass = 37.8 g Mass of mineral isthanks and hope you like the answer